OpenStudy (anonymous):

So when you are trying to find the limit and you know the x values like, 0.5, 0.9, 0.99, 0.999,0.9999, etc. How do you find the f(x) values so that you can find the limit?

jimthompson5910 (jim_thompson5910):

what is the f(x) function?

OpenStudy (anonymous):

(x2-1)/(x-1) as x approaches 1

jimthompson5910 (jim_thompson5910):

oh you can just factor x^2 - 1 to get (x-1)(x+1) notice how the x-1 terms cancel

jimthompson5910 (jim_thompson5910):

after cancellations, you're left with x+1 then you plug in x = 1 to get x+1 = 1+1 = 2 so the limit is 2

jimthompson5910 (jim_thompson5910):

or you can plug in values of x into the function that get closer and closer to x = 1 you'll find that y will get closer and closer to 2